OpenStudy (anonymous):

if f is the inverse of g, then we know that f(g(x))=x use this fact to derive the derivative formula dy/dx e^x=e^x

OpenStudy (anonymous):

is it \[\frac{dy}{dx}e^{x} = e^{x}\] or \[\frac{d e^{x}}{dx} = e^{x}\]

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

its (e^x)' = e^x